OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative.

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@zepdrix @Kainui

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Ok, did you find the derivative, or that is the problem you are having?

OpenStudy (solomonzelman):

In general, \(\color{#000000}{ \displaystyle \frac{d}{dx} (a)=0 }\) \(\color{#000000}{ \displaystyle \frac{d}{dx} (ax)=a }\)

OpenStudy (anonymous):

i need help on the whole question..

OpenStudy (solomonzelman):

Ok, do you know what the «derivative» really means conceptually?

OpenStudy (solomonzelman):

I mean, do you know that the derivative is the slope, or is it just something that you have once heard... ?

OpenStudy (anonymous):

okay so according to the lesson, i find the derivative by using the difference quotient

OpenStudy (solomonzelman):